### v1

# Hiba Arnaout1, Noy Sternlicht2, Tom Hope2,3, Iryna Gurevych1

Abstract
Understanding the impact of scientific publications is crucial for identifying breakthroughs and guiding future research. Traditional metrics based on citation counts often miss the nuanced ways a paper contributes to its field. In this work, we propose a new task: generating nuanced, expressive, and time-aware impact summaries that capture both praise (confirmation citations) and critique (correction citations) through the evolution of fine-grained citation intents. We introduce an evaluation framework tailored to this task, showing moderate to strong human correlation on subjective metrics such as insightfulness. Expert feedback from professors reveals a strong interest in these summaries and suggests future improvements. Data and code are made available.  [^1] https://ukplab.github.io/arxiv2025-generating-impact-summaries
In-depth Research Impact Summarization
through Fine-Grained Temporal Citation Analysis
Hiba Arnaout1, Noy Sternlicht2, Tom Hope2,3, Iryna Gurevych1 1UKP Lab, TU Darmstadt and Hessian Center for AI (hessian.AI) 2School of Computer Science and Engineering, Hebrew University of Jerusalem 3The Allen Institute for AI (AI2) www.ukp.tu-darmstadt.de
1 Introduction
Figure 1: We propose a new task to summarize a paper’s evolving impact over time, by analyzing impact-revealing citation contexts, reflecting both praise (confirming ideas) and critique (calls for correction). Our summary of  Koller and Sahami (1997) reveals its impact trajectory—adaptation, critique, and rediscovery—offering deeper insight than its $\sim$1.4k cite count.
Citation counts are a common proxy for measuring the impact of research papers, but they offer only a shallow view that fails to capture how a paper has influenced subsequent work. A raw citation count does not reveal whether a paper was foundational, extended, critiqued, or merely mentioned in passing. To truly understand the impact of a research paper, we must go beyond simple counts and examine the context in which it is cited, analyzing how its ideas have been discussed, applied, and evolved over time. Manually tracking how a paper is discussed across a very large number of publications and diverse domains is infeasible due to the scale and complexity of the task. To address this, we introduce a new task: research impact summary generation, which aims to generate concise, time-aware narratives that reflect a paper’s evolving scientific influence. These summaries can support practical use cases such as enhancing future research directions by helping researchers quickly assess a paper’s relevance and legacy, or assisting funding agencies, hiring committees, and research administrators in making more informed evaluations of scholarly contributions. Figure 1 shows an example of an impact summary about a research paper Koller and Sahami (1997), published at The International Conference on Machine Learning in 1997. This paper’s impact has followed a dynamic trajectory, initially used for its method (1997–2007), later critiqued for scalability and accuracy limitations, then refined (2008–2018), and recently rediscovered as an inspiration for modern methods (2019–present). This example shows the need to look beyond the paper’s $\sim$1.4k citations to understand the deeper story of how ideas gain relevance, face scrutiny, and find renewed significance.
We break the task of generating impact summaries for research papers (Section 2.1) into two subtasks. First, given all the citation contexts of a paper, we identify the “impact revealing” citations that directly interact with the paper, and their specific intents, whether in confirmation of its ideas (e.g., “method use”) or in correction (e.g., “identifying limitations and proposing refinements”). Second, using only the impact-revealing citation contexts, their dates and intents, we generate an impact summary that captures the paper’s influence over time. To the best of our knowledge, this is the first attempt to express scientific impact through time-aware textual summaries derived from citation context analysis. An overview is shown in Figure 1. To implement these subtasks, our method, in Section 2, uses in-context learning (ICL) with Large Language Models (LLMs) to generate fine-grained citation intents in the first stage. In the second stage, it filters for impact-revealing citation contexts based on the generated intents and feeds them to an LLM to generate semi-structured impact summaries under a pre-defined schema.
Existing work on scientific impact analysis predominantly focuses on citation and similar numerical counts as the main indicator of impact Gu and Krenn (2024), ignoring the exact reasons a paper has been cited, which can often be “unimpactful”, e.g., citing a paper to acknowledge background work on a topic. Work on citation context analysis Lauscher et al. (2022) studies the role citations play, but is based on coarse categories and is not directly intended for impact analysis. The goal of our new task is to create a narrative about a paper’s impact using fine-grained citation intent analysis. Moreover, existing work typically emphasizes praise (confirmatory citations)  Zhang et al. (2024); Valenzuela et al. (2015), but scientific progress relies on both confirmation and correction Catalini et al. (2015), as ideas advance through adoption as well as critique and refinement. In this work, we ensure that our generated impact summaries capture both aspects of impact (when applicable), as shown in Figure 1. To support this broader view of impact, we extend the existing PST-Bench data set Zhang et al. (2024), which only focuses on “motivation” and “inspiration” intentions to consider an impact revealing citation, by introducing, in Section 2.5, a new ground-truth data set of 4000 citation contexts labeled for whether they reveal impact. Our extended dataset also covers not only additional intents of confirmatory citations (e.g., method use), but also correction citations (e.g., method refinement). For detailed comparisons with related work, see Appendix A and Section 5.
Finally, in the absence of gold-standard summaries for this novel impact generation task, we propose an automated evaluation framework to measure the trustworthiness and informativeness of our impact summaries (Section 2.4) and demonstrate a moderate to strong correlation with human assessments (Section 3.3). To further validate our approach, we collect expert feedback from professors on the usefulness of impact summaries about their own papers (Section 4), demonstrating the need for such summaries. For future research, we release the code and data used to develop this work.
2 Method
2.1 Task definition
We propose the task of generating time-aware impact summaries for research papers by identifying and analyzing impact-revealing citation contexts and their intents over time.
Definition 1.
Citation context refers to the specific part of a paper $p^{\prime}$ where another paper $p$ is mentioned, including the surrounding text that explains how $p$ is relevant to $p^{\prime}$.
Definition 2.
Citation intent refers to the purpose (the “why”) behind citing paper $p$ in paper $p^{\prime}$, normally inferred from the citation context.
Definition 3.
An Impact-revealing citation intent is a citation intent where the cited paper $p$ has a direct influence on the citing paper $p^{\prime}$ through: (i) Confirmation, by adopting or building upon $p$; or (ii) Critique, by correcting or refining $p$. Citation intents that do not fall under these categories are considered non-impact-revealing.
Figure 1 shows the intent “acknowledge background work”, for the context that simply refers to prior work in a general manner, without actively building upon or challenging it. Hence, this is not considered impact-revealing. In contrast, the intent “method used”, for the context describing the implementation of a popular approach is indeed impact-revealing.
Definition 4.
A scientific impact summary of a research paper describes the impact paper $p$ has had on subsequent research. In other words, how it has been directly used for both confirmation and critique. The summary is time-aware, taking into account the date of these citations, to track the evolution of its impact over time.
A sample impact summary is in Figure 1, complete examples are in Appendix G.5.
2.2 Identifying impact-revealing citation intents
We use ICL with LLMs to generate fine-grained intents and classify them as either “impact-revealing” or “other”. To facilitate this, we manually craft a prompt that incorporates a set of examples covering various citation intents, and ensuring the model can learn from them to generate fine-grained intents and classify them correctly (prompt and examples are in Appendix B and  E.1, respectively). This prompt is designed to capture both the confirmatory and correction citations, allowing for a nuanced understanding of how the cited paper influences the citing paper. It aims to generate concise but expressive labels that clearly explain the reason behind the citation without being overly general or overly specific, e.g., “identifying knowledge gap in literature” (more examples in Appendix E.1). To examine the quality of our method, we conduct experiments in Sections 3.1 and  3.2.
2.3 Generating scientific impact summaries
To generate scientific impact summaries, we begin by identifying impact-revealing citation contexts through the generated fine-grained intents (as detailed in Section 2.2). To ensure the generation of consistent and easily comparable summaries by LLMs, we design a tailored prompt (see Appendix C) that incorporates our definition of scientific impact (see Definition 4). This prompt includes the impact-revealing citation contexts of a given paper, augmented with corresponding fine-grained intents and citation timestamps (i.e., publication years). We evaluate the effectiveness of this approach through an ablation study (Section 3.3) and an expert review (Section 4).
2.4 Evaluation metrics
A key challenge in evaluating our summaries is the lack of gold-standard datasets, due to the task’s novelty. We attempted to construct gold summaries using sources like Wikipedia articles on influential papers and Test of Time Award webpages, but these proved inadequate. Such sources cover only a few papers and offer details about the content of the paper itself [^2] E.g., Wikipedia article for Vaswani et al. (2017): https://en.wikipedia.org/wiki/Attention_Is_All_You_Need. Similarly, Test of Time award announcements [^3] E.g., WSDM 2022 Test of Time Award Winner:https://www.wsdm-conference.org/2022/timetable/event/wsdm-awards-program-test-of-time-presentation/ focus on broad praise, citation counts, and author accomplishments, lacking usage-based insights. Web searches using queries like “what is the scientific impact of paper X” mostly returned different versions of the paper itself, or occasionally, blog posts highlighting novelty, or explaining the paper in simpler terms. Given these limitations, we introduce a new, automated, reference-free evaluation framework with metrics capturing both the trustworthiness and informativeness of the summaries, and later demonstrate its moderate to strong correlation with human judgment (Appendix G.3).
Intent classifier P R F1 Acc
baseline=random 0.54 0.51 0.52 0.50
baseline=always-impact-revealing 0.53 1.0 0.69 0.53
[…]
Meaningful Citations Valenzuela et al. (2015) 0.72 0.46 0.56 0.62
Multi-cite Lauscher et al. (2022) 0.59 0.41 0.48 0.53
Ours 0.74 0.65 0.69 0.69
Table 1: Results for intent classification (“impact-revealing” or “other”). Our classifier is the best at distinguishing between citations that help in understanding the impact of a paper, and those that serve a more general purpose, such as background references or standard acknowledgments.
Trustworthiness. Measuring summary correctness is challenging due to the large gap between input and output lengths, compressing thousands of citation contexts into a few sentences. Without a gold reference, determining what content should be included or excluded becomes difficult. Additionally, time-augmented impact descriptions must be accurate both in content and assigned time period. Inspired by evaluation frameworks in the RAG setting Ru et al. (2024); Shahul et al. (2023); Asai et al. (2024), which aligns with our query-based summarization approach, we define:
(1) Faithfulness: Our time-Aware faithfulness metric examines whether details in the impact summary, i.e., the impact description elaborating on the dominant intent of an impact period, is grounded in the paper’s citation contexts (i.e., the LLM’s input context). We first split the impact summaries into impact descriptions representing different time periods (as defined in the output schema in Appendix C). Next, we instruct an LLM to verify each against citations from the corresponding period. Note that a summary impact description is verified against many citations at once and not in a simple pairwise entailment, since a single impact description can encompass information from multiple citations or discuss trends, which are only possible to verify by inspecting many sources (e.g., “the paper is frequently used for…”). The evaluator LLM is instructed to assess the impact description against the provided citations, determine its faithfulness, and justify the decision with a list of supporting citations. This procedure resembles the task of a human annotator evaluating the faithfulness of a machine-generated text Kim et al. (2024).
(2) Coverage: Coverage is defined as the ratio of impact-revealing intents mentioned in the summary. For example, if a paper is cited for its method use, for inspiring new research, and for exposing limitations and refinement, the summary should reflect all these intents, including relevant details. Inspired by the evaluation rubrics for coverage proposed in Asai et al. (2024), which defines it as the topics mentioned (i.e., the diversity of impact-revealing intents in our case), we develop a two-step evaluation. In the first step, we cluster the list of fine-grained intents under similar topics, e.g., “method used in the legal domain”. These labels might have slightly different granularity, which depends on how much details the members (i.e., intents) of that cluster offer. Next, the evaluator LLM takes both the impact summary and the list of cluster titles that this summary should ideally cover, and returns a list of topics that were actually covered. Finally, we divide the number of intents mentioned in the summary by the total number of impact-revealing intent clusters. Prompts for faithfulness and coverage are in Appendix D.
(3) Citation Year Compliance: We implement this metric as a script that flags citations falling outside the target impact period.
Informativeness. We also assess the informativeness of impact summaries, as trustworthy ones may still be unhelpful if they lack meaningful insights. Grounded in our definitions (Section 2.1), the informativeness of an impact summary depends on its ability to clearly describe the paper’s direct influence on other papers, track the types of influence (intents) over time, and provide details inferred from the citation contexts. We define the following metrics:
(1) Insightfulness measures how well the impact summary articulates the paper’s direct influence on other works.
(2) Trend Awareness evaluates the extent to which the impact summary identifies and distinguishes between different periods of the paper’s influence over time.
(3) Specificity assesses whether the impact summary includes concrete examples, such as techniques, influenced by the paper. To systematically evaluate these criteria, we leverage G-Eval Liu et al. (2023c), a framework that uses LLMs with chain-of-thought (CoT) reasoning to assess the impact summaries. By integrating structured evaluation steps, G-Eval enables LLMs to reason step-by-step, improving the reliability and depth of assessments. We transform our criteria into prompts in Appendix D.
2.5 A new dataset for identifying impact-revealing citation contexts
To construct a new dataset for classifying citation contexts as “impact-revealing” or not, we start by building on the PST-Bench dataset Zhang et al. (2024), which focuses on positive influence through “inspiration” and “motivation”. We augment this by first manually crafting textual patterns for both confirmation and correction intents, and use GPT-4 to generate variations. After crawling 200k random citation contexts (from the Semantic Scholar Academic Graph (S2AG) [^4] https://www.semanticscholar.org/product/api) and checking if they match any of the textual patterns, we randomly sample 2k instances of impact-revealing contexts (50% for confirmation and correction citations each). Additionally, we sample 2k non-impact-revealing citations (“other”) directly from the PST-Bench dataset (references that were not annotated as influential). We also ensure none of these match any of our impact-revealing citation patterns. This results in a balanced 4k citation context dataset, which we release alongside this work. Details on the construction and textual patterns are in Appendix F.1.
3 Experiments
3.1 In-context-learning for fine-grained intent generation
Zero-shot vs. ICL. To evaluate LLMs’ ability to generate fine-grained intents and classify them as “impact-revealing” or “non-impact-revealing”, we manually select 10 citation contexts (Appendix E.1). Our selection criteria was based on diversity, a set that contains impact-revealing (praise, critique) and non-impact-revealing (“other”) intents. The “other” class includes incidental mentions or contexts lacking sufficient information to understand the reason behind it. We use our prompt (Appendix B) with and without these examples to generate and classify intents, using GPT4o-mini, of 200 randomly sampled citation contexts. Two annotators (a postdoc and a PhD student) assess: (i) whether the classification as impact-revealing is correct, (ii) whether the generated intent accurately reflects the citation reason, and (iii) whether the intent is concise, as LLM results can be highly verbose.
Results, including per-field metrics and qualitative examples, are in Appendix E.2. We find that including examples significantly improves performance across all metrics, with gains of 29%, 55%, and 39% in precision, recall, and F1, respectively. Inter-annotator agreement (F1 agreement) is 0.86 for the impact-revealing classifier, 0.77 and 0.85 for the correctness and conciseness of the generated fine-grained intents, respectively.
Effect of different number of shots (K). We conduct an experiment to test our ICL prompt with different Ks. We manually polished the 200 instances from the previous experiment and randomly split the set into train 40%, dev 30%, test 30%. For the number of shots, we test from K=1 to 80. For every dev and test instance, we run the prompt 5 times and report the average, using GPT4o-mini, shuffling the order and selection of the shots. Results are plotted in Figure 2. Increasing the number of shots has an effect on all metrics, outperforming both the ZS (horizontal dashed lines), and the always predict impact-revealing baselines (dotted lines). Performance improves sharply up to K = 50, after which gains are minimal. At K = 50, metrics plateau with precision at 0.90, recall at 0.94, F1 at 0.92, and accuracy at 0.92.
Figure 2: Effect of different number of shots (K) on detecting impact-revealing citations. Zero-shot: horizontal dashed lines; always predict impact-revealing baseline: dotted lines. This shows that adding a reasonable number of examples (around 50) is sufficient to yield a 60% improvement in recall.
Cited Papers Citation Intents (%)
PS MD CS
All 6535 5347 3466
Recent 6535 5743 5842
[…]
Highly cited 6634 3664 3862
[…]
Table 2: Recent = the last 5 years, Highly cited = top 20 % by citation count in our dataset; orange for impact-revealing, light blue for other; PS= psychology, MD = medicine, CS = computer science.
How does citation intent vary across fields? Table 2 analyzes citation intents across 70k citation contexts from psychology, medicine, and computer science papers. It shows that, generally, psychology citations tend to lean towards impact-revealing, computer science towards the non-impact-revealing (“other”) side, with the exception of more recent papers, and medicine has a more balanced split. Looking at specific examples, we notice that citations in psychology papers often have a stronger subjective tone, particularly in correction citations. Moreover, in recent years, computer science shows a shift toward more impact-revealing citations, likely reflecting the novelty and immediate impact of AI research before entering the "legacy" phase with more background-oriented citations. Details and examples are in Appendix F.
Prompt variant Trustworthiness Informativeness
Citations Intents Faith. Cov. Cov.@3 Cyc. Insi. Trend. Spec.
None 0.77 0.25 0.58 n/a 0.70 0.94 0.75
All 0.83 0.32 0.74 0.55 0.80 0.95 0.85
All 0.84 0.32 0.73 0.48 0.80 0.97 0.86
Impact-rev. 0.87 0.33 0.73 0.59 0.80 0.96 0.87
Impact-rev. 0.88 0.34 0.75 0.56 0.83 0.98 0.88
Table 3: Ablation results - Faithfulness: Faith., Coverage: Cov., Citation Year Compliance: Cyc., Specificity: Spec., Insightfulness: Insi., Trend Awareness: Trend. These results show that adding impact-revealing contexts and their intents has an improvement on almost all metrics.
3.2 Can existing intent classifiers detect impact-revealing citations?
Comparison with existing methods. We compare our impact-revealing classifier against the following popular intent classification methods: (i) Meaningful Citations Valenzuela et al. (2015): a supervised machine learning method to classify citation intents into meaningful or non-meaningful, by leveraging context and citation metadata; (ii) Structural Scaffolds Cohan et al. (2019): a multi-task model which enhances the main task by integrating auxiliary tasks; (iii) Multi-cite Lauscher et al. (2022): a multi-label classification method that predicts multiple intents using a fine-tuned SciBERT Beltagy et al. (2019). We map their intent classes to either “impact-revealing” or “other”. More details are provided in Appendix F.2. For our method, we use the ICL prompt, with K=50 and LLM=GPT4o-mini. For every test instance, we run our prompt 3 times (shuffling the order of the shots), and take a majority vote for the classification. We note that the results of the 3 runs have a full agreement of 72%.
Results. Numerical results are reported in Table 1. Our method demonstrates the best performance on this task, with the most significant improvement in recall, outperforming the second-best method by 19 percentage points. This is particularly crucial for our goal in generating impact summaries from impact-revealing citations because the highest recall in this context means that our method ensures that fewer meaningful citations are missed. The qualitative results are shown in the appendix F.2.
3.3 Ablation study for impact summary generation
Data: papers, citations, intents. We select 105 papers from the Semantic Scholar Academic Graph (S2AG) [^5] https://www.semanticscholar.org/product/api, from psychology, medicine and computer science (35 each), published between 1974 and 2022 at top venues in their respective fields. The citation count for each article ranges from $\sim$ 500 to $\sim$ 5000. We collect citations and their contexts using the citation lookup feature in the S2AG API. We first run our fine-grained intent generator (with K=50) and classifier for every citation context. This amounts to 70k citation contexts with their generated intents and their classes.
Prompt variants. We experiment with different input configurations to assess their effect on the summary’s quality, focusing on: (1) Citation contexts, including all, only impact-revealing, or none, to test what the LLM infers without citations and how varying context types affect the summary. Although full citation contexts can enrich summaries, they can also introduce noise from incidental references Liu et al. (2023a). (2) Fine-grained intents, whether to include the intents alongside citation contexts.
Table 3 lists all variants. In total, we generate 945 impact summaries6661 variant without citations + (4 citation-based variants * 2 orderings) * 105 papers. See Appendix C for the prompt and schema. For the intent generation task, we choose GPT-4o-mini due to its lower cost. For both impact generation and evaluation, the more advanced tasks, we use GPT-4o.
Results. Table 3 shows the results for all metrics. Even though it shows a good understanding of how the paper was cited, likely due to the LLM’s training data having access to a large corpus of scholarly articles, the zero-shot variant is the least faithful, and we observe that prompts with only impact-revealing citations generate more faithful impact summaries. This suggests that longer contexts might induce hallucinations in the generation process. Interestingly, providing the citation intents also gives a slight boost to the faithfulness of the impact summaries. A possible reason is that the intents encourage wording that aligns with the input paper’s citations.
Coverage is especially challenging, requiring a balance of completeness, relevance, and conciseness. Still, our variant outperforms the baseline by 9%, showing the LLM can capture key themes. Some inputs contain over 100 distinct themes, explaining modest overall scores. When evaluating only the top-3 most frequent themes (the largest 3 clusters), coverage rises significantly, as our best variant reaches 0.75 (+17%), highlighting its ability to capture the most important aspects of a paper’s impact. Variants exposed only to impact-revealing citations show higher citation year compliance, while adding citation intents lowers it, likely because models focus more on intent than timing. Examples are in Appendix G.1, with further discussion in Section 7.
All variants score high on trend awareness, with a slight edge for the impact-revealing + intents variant, showing LLMs can recognize distinct impact periods. This variant also improves insightfulness and specificity by 13% over the zero-shot baseline. Qualitative examples are in Appendix G.1. G-Eval, via DeepEval [^7] https://github.com/confident-ai/deepeval, also offers reasons for each score (examples in Appendix G.4). We visualize sample impact summaries in Appendix G.5. To examine whether results change depending on the field of the paper, we report results per field in Appendix G.2.
LLM-Human correlation We compute F1 agreement between human and LLM evaluators (50 instances per metric), yielding scores of 0.83 for faithfulness, 0.71 for coverage, 0.66 for insightfulness, 0.86 for trend awareness, and 0.74 for specificity. Correlation metrics (Spearman and Kendall-Tau) in Appendix G.3 show moderate to strong statistical significance.
3.4 Practical applications
We showcase the usability of our work in two practical applications. First, by analyzing 10 diverse research topics, such as LLMs for code generation, open information extraction, commonsense knowledge mining, and comparing papers with similar citation counts, it shows that papers can have very different types of impact. For example, in open information extraction, some papers are cited to highlight limitations or motivate new work Cui et al. (2018), while others are cited mainly for their methods Gashteovski et al. (2017); Han et al. (2019). More details and examples are in Appendix I.1.
Second, we explore generating author-level impact summaries by aggregating LLM-generated paper-level summaries for an author’s top-cited papers. Examples for two senior NLP researchers are shown in Appendix I.2. This approach has strong potential for broader applications at the level of labs, institutions, or venues, which we leave for future work.
Figure 3: Pairwise comparison. Relevance: Which summary better reflects the paper’s actual impact? Insightfulness: Which summary offers more valuable or novel information about the paper’s usage? Professors consider our summaries more relevant and insightful.
4 Human Evaluation
Setup. To gain deeper insights into the quality of our impact summaries and collect expert feedback for potential future improvements, we conduct a user study with 9 university professors from diverse backgrounds (details in Appendix H.1). Each expert reviews impact summaries of their own papers through a two-part evaluation: (1) a pairwise comparison, where they choose the better summary based on relevance and insightfulness, comparing the zero-shot variant against our best variant. To prevent bias, we alternate the positions of the impact summaries. (2) a perceived usefulness assessment, where experts rate their agreement with given statements about an impact summary (generated by our best variant), using a 1-5 Likert scale. The professors evaluated the impact summaries for 82 papers and offered open-ended feedback.
Results. Figure 3 shows results of the pairwise comparison. Impact summaries generated by our variant [^8] Impact-revealing citations with intents. demonstrate greater relevance (63%) and insightfulness (75%) than the baseline [^9] Zero-shot variant.
Results of the perceived usefulness (visualized in Appendix H.2) show that approximately 60% of professors agreed that the summaries provided an appropriate level of detail (clarity) and offered novel insights into how their papers were used, i.e., information not readily available elsewhere. Notably, for papers in the top 10% based on impact-revealing citations, agreement on clarity and informativeness rose to 75%. Results and a discussion on open feedback are in Appendix H.3.
5 Related Work
Scientific impact analysis. Studying scientific impact is vital for guiding research and recognizing contributions. While prior work relies on citation counts and related metrics Hutchins et al. (2016); Bos and Nitza (2019); Siudem et al. (2020); Min et al. (2021); Wahle et al. (2023); Gu and Krenn (2024), impact cannot be reduced to a single number Zhu et al. (2015). Few studies use citation context to assess scientific impact, and those that do focus on classifying citation intent individually rather than aggregating them in summaries Jurgens et al. (2018); Valenzuela et al. (2015). Our approach uses fine-grained citation context analysis to capture and aggregate specific citation reasons across impact-revealing citations.
Other work links impact to novelty of the paper’s content Rüdiger et al. (2021); Shi and Evans (2023); Arts et al. (2021), but overlooks how papers are used. Most also equate impact with praise Zhang et al. (2024); Valenzuela et al. (2015); Zhu et al. (2015), whereas we distinguish confirmatory from correction citations, offering a fuller view. As detailed in Appendix A, we are the first to generate time-aware, multifaceted impact summaries from fine-grained citation analysis.
Citation intent prediction. Prior work on citation intent classification uses pre-defined categories (e.g., method use, background) with early methods relying on supervised learning and linguistic features Teufel et al. (2006); Jurgens et al. (2018); Tuarob et al. (2019). Later approaches applied multi-task learning Cohan et al. (2019), while Lauscher et al. (2022) introduced a multi-label method using fine-tuned SciBERT Beltagy et al. (2019). These rely on coarse categories; in contrast, we use in-context learning to generate fine-grained, free-form intent descriptions (Table F.2), offering more precise insights, an approach recently shown to aid related work generation Şahinuç et al. (2024).
Multi-document summarization in the era of LLMs. Recent work in query-focused multi-document summarization uses LLMs to generate context-aware summaries Roy and Kundu (2023), with approaches like graph-based QA over private corpora Edge et al. (2024) and retrieval-augmented summarization Zhang et al. (2025). In the scientific domain, prior efforts focus on summarizing papers for related work Şahinuç et al. (2024); Lu et al. (2020); Chen et al. (2021); Wu et al. (2021) or reviews Kasanishi et al. (2023). In contrast, we generate free-text summaries of scientific impact using a novel time-aware, multi-document approach.
LLM Evaluators. Retrieval-Augmented Generation (RAG) systems condition generation on external information retrieved from user queries Yu et al. (2024); Gao et al. (2023), but evaluating them is challenging due to multiple stages like chunking and retrieval. Tools like RagChecker Ru et al. (2024) use entailment-based metrics to assess faithfulness, while RAGAs Shahul et al. (2023) offers reference-free evaluation of relevance and faithfulness. Inspired by these, we assess our impact summaries’ faithfulness and coverage by validating them against the citation context corpus.
The LLM-as-a-judge paradigm, used for tasks like QA and writing evaluation Zheng et al. (2023); Shao et al. (2024); Asai et al. (2024), enables scalable assessment of complex qualities. Given the novelty of scientific impact summarization, we adopt G-Eval Liu et al. (2023b), a CoT-based framework, to assess informativeness.
6 Conclusion
This work introduces a novel approach for generating time-aware impact summaries of research papers by analyzing evolving fine-grained confirmatory and correction citation intents. Our approach goes beyond citation counts, offering a more nuanced understanding of a paper’s impact over time. Automated evaluations align moderately to strongly with human judgments, and expert feedback highlights the value and potential for refinement of our summaries. We release our data and code to support research in this new task.
7 Limitations
Language. Our work focuses on English papers, as it is the dominant language in most research fields. Extending this approach to a cross-lingual setting is a promising future direction.
Human evaluation. Constructing a large-scale user study is challenging because assessors need deep knowledge of the papers’ impact. We focused on experts evaluating their own work, but this limited our pool to 9 experts, as highly cited papers are typically authored by busy professors. We chose these experienced authors with multiple impactful papers to maximize the number of impact summaries evaluated per person.
Trustworthiness evaluation. Without restrictions, our approach shows unsatisfying coverage (0.34) and citation year compliance (0.59), compared to other metrics such as faithfulness. Regarding the year compliance, citation phrases often include multiple references accompanied by publication years, which can confuse the model and lead to incorrect citation year associations. In the future, we plan to remove or distinguish unrelated, potentially distracting numbers from the citation context (e.g., using well-crafted heuristics or prompt-based methods). For coverage, the LLM may focus on certain prominent themes while underrepresenting others, especially with complex topics. We briefly reported the coverage numbers for the most frequent themes, which showed a great improvement. In the future, we plan to explore this further, for instance, by ranking the importance of the themes and factor this into the coverage computation at different ranks. However, for this study, we decided to test full coverage to make sure the model captures a wide range of themes, without focusing on any specific ones, to check its overall understanding of the paper’s impact. Finally, although checking potentially hundreds of citation contexts/intents at once raises concerns, the strong LLM-human agreement reported in Section 3.3, 0.83 F1 agreement for faithfulness and 0.71 for coverage, suggests such issues are unlikely.
Experimenting with more LLMs. In this paper, we prioritize introducing and exploring the new task in depth, which is why we only experimented with the long-context model GPT-4o. We leave testing a wider range of models for future work. Although using the same LLM for both generation and evaluation can introduce bias, it will also promote consistency in task interpretation, and since our evaluation is not preference-based but scoring-based, bias is unlikely to affect results.
Ethics Statement
The data used in this study is publicly accessible and provided under open licenses, ensuring long-term reproducibility and building upon our work. For the expert study, we do not disclose the names of participating professors or publish individual votes and verbatim feedback. Instead, we report aggregated votes and rephrased feedback for suggestions of future improvements. While this study explores a promising new task, we caution that inaccuracies in the generated summaries could mislead readers about a paper’s actual impact. Therefore, we do not recommend using our method directly and without human verification for critical decisions, such as academic hiring or grant allocations.
Appendix A Existing Work on Scientific Impact Analysis
In Table 4, we compare our work with existing research on scientific impact analysis. More on this in Section 5.
Work Impact as Facets Time Method Intents Fields
Valenzuela et al. (2015) Influential-citation counts P. Supervised Machine Learning coarse Computer science
Zhu et al. (2015) Influential-citation counts P. Supervised Machine Learning 10 fields
Hutchins et al. (2016) Relative citation ratio n/a Statistical measures Medicine
He and Chen (2018) Quantified change rate of citation context n/a unsupervised temporal embedding analysis Biomedical
Jurgens et al. (2018) Citation counts n/a Supervised Machine Learning coarse Computer Science
Bos and Nitza (2019) Citation ratio n/a Statistical measures 13 fields
Siudem et al. (2020) Bibliometric indexes n/a Statistical measures Computer science
Rüdiger et al. (2021) Word lists n/a Text clustering Information systems
Arts et al. (2021) Concept combinations P. Text processing U.S. patents
Min et al. (2021) Citation cascades n/a Network analysis Physics
Wahle et al. (2023) Citation Field Diversity Index n/a Statistical measures 23 fields, focus on Computer science
Shi and Evans (2023) Concept combinations P. Hypergraph model Medicine, Physics
Zhang et al. (2024) Influential-citation counts P. LLM Computer science
Gu and Krenn (2024) Citation counts n/a FNN Physics
Arts et al. (2025) Paper counts P. Statistical measures Nobel Prize-winning papers
This work Textual Summary P., C. LLM fine-grained Computer science, Psychology, Medicine
Table 4: Comparison with existing work on scientific impact analysis. Our work is the first to express scientific impact through time-aware textual summaries derived from fine-grained citation context analysis, which covers both praise (confirmation) and critique (correction).
Appendix B Prompt for Identifying Impact-revealing Citation Intents
The prompt used is shown in Figure 4.
Identifying impact-revealing intents A citation context in a scientific paper refers to the specific part of a paper $p^{\prime}$ where another paper $p$ is mentioned, including the surrounding sentences or paragraphs that explains how and why $p$ is relevant to $p^{\prime}$. A citation context with an impact-revealing intent is a type of citation in scientific writing that highlights the significance or influence of a previously published work, often emphasizing its contribution or importance to the current research or the broader field, e.g., its role in inspiring, motivating, supporting, filling gaps, critically analyzing, or contributing methods, tools, data, extensions, or benchmarks for the current research. Other types of intents include a reference to prior work in a scientific paper that provides background or context without emphasizing the impact, significance, or influence of the cited work. It acknowledges the source in a routine or supporting role rather than showcasing its importance to the research. Given a citation context, describe, in a few words, the intention behind this citation phrase. Then, decide on the category of this intention. In particular, whether the intention behind this citation phrase is impact-revealing or not (i.e., incidental or that there isn’t enough information to realize the real intention behind it). For the intention category, only return one of the following two labels impact-revealing or other. Below are examples: \$examples\$
Figure 4: Prompt for generating and classifying fine-grained intents.
Appendix C Prompt for Generating Impact Summaries and Output Schema
To ensure that the generated output contain all the components required to construct the impact summaries, (namely the impact periods, their dominating intent, details about the period, and citations used as evidence), we make use of OpenAI’s structured output feature [^10] https://platform.openai.com/docs/guides/structured-outputs. Moreover, this facilitates a more systematic automated evaluation for our prompt variants. The prompt design is illustrated in Figure 5, and the corresponding structured output schema is presented in Figure 6.
Generating an impact summary about a research paper The scientific impact summary of a research paper describes the impact a given paper had on other papers, including both praise and critique. To understand the impact of a paper, one needs to understand how exactly it has been utilized and discussed by other papers. This is normally referred to as citation intents. One also needs to understand the evolution of the impact and citation intents over time. Given an input paper’s title, its publication year, and its citation context, describe the impact of that paper. The citation context includes five components: $<$citation ID, citation title, citation year, citation context, citation intent$>$. Given the input paper with id \$paperId\$ titled \$title\$ published in \$year\$, and the following list of papers citing it: \$citation_context\$ Generate an impact summary about the input paper.
Figure 5: Prompt for generating an impact summary about a research paper.
1# Schema for structured output
2{"name": "impact_statement",
3"schema": {"type": "object",
4 "properties": {
5 "input_paper_info": {
6 "type": "object",
7 "items": {
8 "input_paper_id": "id of input paper",
9 "input_paper_title": "title of input paper",
10 "input_paper_year": "year of input paper"}},
11 "impact_periods":{
12 "type": "array",
13 "items": {
14 "impact_period": "start year - end year",
15 "aspect_of_period": "the dominating citation intent(s) of that period",
16 "impact_description": "a paragraph to describe the impact of that period",
17 "evidence": "citing papers from that period to back up the described impact aspect"}}
18 }
19 }
20}
Figure 6: Output schema for impact summary generation.
Appendix D Prompts for Evaluating Impact Summaries
Trustworthiness metrics: The prompt we use for our faithfulness evaluation is shown in Figure 7, and for our two-step coverage evaluation in Figure 8. Informativeness metrics: The evaluation steps are shown in Figure 9.
Faithfulness evaluation prompt **Task:** Verify the faithfulness of an impact description regarding the paper "{{PAPER_NAME}}". It is faithful if it can be supported by one or more of the paper’s citations from the provided list. Each citation is formatted as <title>:citation_text, where the title indicates the title of the citing paper.
**Impact description to Verify:**
<impact-description>
{{DESCRIPTION}}
</impact-description>
**Citation List:**
<citations>
{{SOURCES}}
</citations>
**Steps to Complete the Task:**
1. Understand the impact description and its specified time period.
2. Review each citation in the provided list.
3. Determine if the impact description can be supported by any single citation or a combination of citations.
4. If the impact description is supported, identify the relevant citations that support it.
5. If the impact description cannot be supported or is contradicted by the citations, determine it as unfaithful.
**Response Format:**
- **Analysis:** <analysis> [Provide your analysis here] </analysis>
- **Answer:** <answer> [yes/no] </answer>
- **Proof:** <proof> [List the exact text of the citations that support the impact description, or "none" if unfaithful] </proof>
**Additional Guidelines:**
- The answer should be "yes" or "no" only.
- In the proof section, include only the exact text of relevant citations without explanations.
- List all necessary citations if multiple are needed to support the impact description.
- If the impact description is unfaithful, state "none" in the proof section.
- Avoid additional commentary outside the specified format.
Figure 7: Prompt for verifying the time-aware faithfulness of scientific impact summaries.
Prompt for coverage evaluation Step 1: Given this list of phrases \$listOfPhrases\$, cluster highly similar phrases and give a label (expressive theme) for every cluster. Step 2: Given this list of themes in the format of a python list: \$listOfThemes\$, determine how many of these themes were implicitly or explicitly mentioned in this summary \$summary\$, and list them.
Figure 8: Prompt for measuring coverage of an impact summary.
1from deepeval.test_case import LLMTestCase, LLMTestCaseParams
2from deepeval.metrics import GEval
3
4insight_metric = GEval(
5 name="Insightfulness",
6 evaluation_steps=[
7 "Determine whether the impact summary describes how the paper has been directly used by or influenced by other works.",
8 "Assess how well the impact summary articulates the paper’s influence with informative details.",
9 "You should heavily penalize the impact summary for lack of insight."
10 ]
11)
12trend_metric = GEval(
13 name="Trend awareness",
14 evaluation_steps=[
15 "Determine whether the impact summary mentions how the impact of the paper has changed over time, ensuring each impact period is clearly identified with descriptive titles.",
16 "You should heavily penalize the impact summary if the titles of consecutive impact periods are not diverse."
17 ]
18)
19specif_metric = GEval(
20 name="Specificity",
21 evaluation_steps=[
22 "Determine whether the impact summary mentions specific techniques, frameworks, or studies influenced by the paper, or if it remains broad and lacking supporting details.",
23 "You should heavily penalize the impact summary if it only restates the title and abstract or provides vague, generic statements without concrete examples of the influence of the paper."
24 ]
25)
Figure 9: Code snippet for chain of thoughts (CoTs) evaluation steps for 3 metrics of the impact summary evaluation.
Appendix E Details about Zero-shot vs. ICL for Fine-grained Intent Generation
E.1 Training examples
The list of manually annotated examples used in this experiment are in Table 5.
Citation context Intent Class
In order to reduce the memory requirements, we apply a minimization process [1]. use of minimization methodology impact-revealing
Motor adaptation is the process of re-shaping acquired motor skills through the reduction of errors (Hardwick and Celnik, 2014; Krakauer, 2009) defines the term motor adaptation other
Moreover, none of the above studies explored whether temporally primary PLEs are associated with an increased risk of subsequent insomnia [1,2].. In this study, we explored the changes in prevalence of insomnia and PLEs before and during the pandemic. identifying and addressing knowledge gap in literature impact-revealing
Chiu and Nichols (2016) introduced convolutional neural networks for NER background about NER methods other
We employ the single-link method to compute the similarity be-tween two clusters, which has been applied widely in prior research (Bagga and Baldwin (1998); Mann and Yarowsky (2003)) use of cluster similarity methods impact-revealing
Quirk and Poon (2017) and Peng et al. (2017) build two distantly supervised datasets without human annotation, which may make the evaluation less reliable. In this paper, we present DocRED, a large-scale human-annotated document-level RE dataset.. criticizing existing datasets and proposing a better one impact-revealing
In adults with severe malaria, increased Ang-2 plasma levels were associated with a decrease in NO bioavailability, higher lactate plasma concentrations, and patient mortality [101]. reporting on existing studies about malaria other
, similarity measures [3]) to select an answer. not enough information other
Inspired by the reference [4], we proposed a novel algorithm called Soft-DDQN and applied it to the robot PAP skill learning problem drawing inspiration from prior work to propose a new algorithm impact-revealing
For instance, while Yoon et al. (2018) show that expanding the network capacity thorough width is helpful, they have not studied the impact of increasing capacity when the depth increases, nor why increasing the width is helpful. Overall, in this work, we are interested in understanding the impacts of network structure (e.g., width and depth).. highlighting gaps in existing research on network capacity impact-revealing
Table 5: Training examples (manually created from real contexts) for fine-grained intent generation and classification.
E.2 More on setup and results
We use GPT-4o mini with temperature 0. As for the source of papers and citation context, we use the Semantic Scholar Academic Graph (S2AG) [^11] https://www.semanticscholar.org/product/api.
Figure 10 shows the quantitative results per paper field, and Table 7 shows qualitative examples. We do not observe any large variations indicating that our method is not limited to a specific field. However, the highest recall (a perfect score of 1.0) was achieved in the psychology field. This may suggest that citation contexts in psychology papers tend to express intentions more explicitly, particularly those that reveal impact, e.g. “Some of those assumptions have been controversial, as researchers disagree about whether the kinds of behaviors measured by particular implicit tests should be considered indicators of attitudes or something else”. This seems to be in line with our observations made during error analysis. To understand where our ICL prompt failed, we plot the confusion matrices in Figure 11 and inspect the 10% impact-revealing citation phrases that were misclassified as “other”. We notice that the classifier considers minor “resource use” intent as non-impact revealing, e.g., “We retrieved 400 results for each keyword, resulting in a total of 145,682 stories downloaded in the JavaScript Object Notation (JSON) format [48].” On the same note, it appears that when “inspiration” is only hinted at in the citation context, the classifier is not able to pick up the signal, e.g., “The search problem here corresponds closely to a binary dynamic constraint satisfaction problem [Mittal & Falkenhainer , 1990]”. Inter-annotator agreement on the annotated results is in Table 6. They range from good to strong.
Figure 10: Results for fine-grained intent generations (ZS vs. ICL with 10 examples). Adding examples show significant improvements on metrics like precision, recall, F1.
Task F1 agreement
Impact-revealing classifier 0.86
Fine-grained intent: correctness 0.77
Fine-grained intent: conciseness 0.85
Table 6: Inter-annotator agreement on fine-grained intent generation results (Section 3.1).
Figure 11: Confusion matrix for the impact-revealing classification task. More insights into misclassified instances in Section E.2
Citation context Baseline Ours
Policy optimization is performed using an implementation of TRPO from rllab [62] with a step size of 0. The intention behind this citation phrase is to acknowledge the use of a specific tool or method (TRPO from rllab) in the research without emphasizing its impact or significance other reporting method used for policy optimization impact-revealing
Alternative methods that instead rely on constraining model parameters, so-called regularization approaches [18,37], have in turn been shown to perform poorly on medical data. This citation phrase acknowledges prior work related to regularization approaches and highlights that these methods perform poorly on medical data. other highlighting the limitations of regularization approaches on medical data impact-revealing
Additionally, suffering can occur due to illness, to the way in which illness is experienced, or, even, in absence of illness (Cassell, 1982; Frank, 2001). The citation phrase acknowledges prior work to provide context and background for the concept of suffering in relation to illness and its experience. It doesn’t emphasize the impact or significance of the cited works. other providing context on the nature of suffering other
Table 7: Sample generated intents, baseline= ZS, Ours=ICL (10 examples).
Appendix F Intents across Fields
In Table 2, we show an analysis of citation intents in the psychology, medicine, and computer science fields, each also divided by recency and citation counts. It shows that, generally, psychology citations tend to lean towards impact-revealing, computer science towards the non-impact-revealing (“other”) side, with the exception of more recent papers, and medicine has a more balanced split. Looking at specific examples, we notice that citations in psychology papers lack an objective tone, especially when it comes to critique, e.g., “some of those assumptions have been controversial”, and “researchers disagree about whether the kinds of behaviors measured by particular implicit tests should be considered indicators of attitudes or something else”. However, it does seem that in computer science there is a noticeable shift toward more impact-revealing citations in recent years. This shift likely reflects the novelty of recent research and the immediate impact of AI papers, as these papers have not yet entered the “legacy” phase, i.e., when papers start to receive more citations with “background work”-like intents. It is worth mentioning that these “legacy” papers are naturally strongly present in the highly cited subsets (by citation counts).
F.1 A new dataset for identifying impact-revealing citations
Since our task on classifying citation context into “impact-revealing” (for both confirmation/praise and correction/critique) or “other” is new, there is a lack of groundtruth data. To construct such a resource, we build upon on an existing human annotated dataset, namely PST-Bench Zhang et al. (2024). An instance in this dataset consists of a pair of research papers annotated for whether one of them influenced (impacted) the other. Influence here is defined only through a positive lens, and is further restricted to the following two intents: “inspiration” and “motivation”. To create our dataset, first, we randomly select 1k influential (impact-revealing) papers, as labeled by the PST-Bench dataset annotators, and look up their aggregated citation contexts [^12] The cited paper might be mentioned several times in the citing paper.. Since the meaning of influential, according to PST-Bench, only focuses on two intents “inspiration” and “motivation”, both limited to praise, we augment the dataset with other types for both confirmatory and correction citation intents. To do so, we manually craft a handful of textual patterns such as “has led to open questions/challenges/unresolved issues” to capture intents indicating research limitations, then ask GPT4-o to provide a longer list with variations of these patterns. We end up with a total of 33 patterns (examples shown in Figure 12). Following this step, we crawl $\sim$ 200k aggregated citation contexts and search them using the list of impact-revealing phrases, then randomly sample 1k instances from the matching cases. To augment the data with non-impact-revealing (“other”) examples, we sample 2k non-influential instances from the PST-Bench dataset. These are citations that are listed under the category “other citations”, as in non-influential ones. To ensure the deemed non-influential examples are not in fact impact-revealing (ones which belong to intents other than “motivation” and “inspiration”), we verify that there is no match with all the impact-revealing phrases. Our newly constructed dataset, which we release with this work, consists of 4k citation contexts augmented with either “impact-revealing” or “other” intent class (balanced).
1# Python list containing impact-revealing phrases (both praise and critique)
2impact_revealing_phrases = [
[…]
9 ...
10]
Figure 12: Samples impact-revealing textual patterns (33 in total). These were used to extend the PST-Bench dataset to include more diverse impact-revealing citations (especially critique/correction).
F.2 Details about comparisons with intent classifiers
Mappings of the baselines’ intents to ours are shown in Table 8. Two of the external methods for intent classification, namely the work of  Cohan et al. (2019) and  Lauscher et al. (2022) allow for multi-label classification. This means: a citation context can receive both “impact-revealing” and “other” labels (e.g., with intents “Result” and “Background”). However, in these cases, a ranking of the predictions can be utilized. For  Lauscher et al. (2022), we use the class with the highest prediction probability. For  Cohan et al. (2019), we count the frequency of every intent class describing the concatenated citation context, and most frequent class represented in the context. We break ties with a random pick.
Intent Classifier Classes
Structural Scaffolds Cohan et al. (2019) Background other
Method impact-revealing
Result impact-revealing
Meaningful Citations Valenzuela et al. (2015) non-meaningful other
meaningful impact-revealing
Multi-cite Lauscher et al. (2022) Background other
Motivation impact-revealing
Future Work other
Similar/Difference impact-revealing
Uses impact-revealing
Extension impact-revealing
Table 8: Mapping existing coarse intent classes of existing intent classifiers to our “impact-revealing” or “other” classes.
Sample results are shown in Table F.2.
Citation context: “Inspired by the theory of hierarchical abstract machines (Parr and Russell 1998), we cast the task of profile reviser as a hierarchical Markov Decision Process (MDP).
[…]
Scaffolds
Meaningful
Citations
Multi-cite Ours
[…]
(non-meaningful)
impact-revealing (uses) impact-revealing (drawing on theoretical foundations for task formulation)
Citation context: “..on social network datasets, it is quite intuitive trying to extract information from text data to do ideology-detection, only a few paid attention to links [9, 13]… Even though some realized the importance of links [9, 13], they failed to provide an embedding.
[…]
Scaffolds
Meaningful
Citations
Multi-cite Ours
[…]
other
(non-meaningful)
impact-revealing (motivation) impact-revealing (highlighting the gap in ideology detection methods)
Citation context: “Recurrent networks are dedicated sequence models that maintain a vector of hidden activations that are propagated through time (Elman, 1990; Werbos, 1990; Graves, 2012).
[…]
Scaffolds
Meaningful
Citations
Multi-cite Ours
[…]
other
(non-meaningful)
other
other
(providing context on recurrent networks)
Table 9: Intent classes predicted by existing work on intent classification. The intents between parentheses are the actual labels predicted by the method, later mapped to either “impact-revealing” or “other” based on the mappings shown in Table 8. For our method, the output is both the intent and its class.
Appendix G Details about the Ablation Study
G.1 Examples
Table 10 shows examples of verified (faithful) impact descriptions. Table 11 shows examples for coverage evaluation. Note that we instruct the LLM to return the actual list of topics covered instead of only the number of themes covered, in order to determine its performance compared to a human evaluator (see Table 12). Table 13 shows examples from the citation year compliance evaluation. In Table 14, we show qualitative examples in informativeness.
G.2 Results per field
Table 15 presents the results for papers of different research fields. We observe that faithfulness and citation year compliance exhibit similar patterns in specific fields as in the overall results, with variants with impact-revealing citations achieving better results. The only exception is the computer science, with a minor difference on year compliance of 1% in favor of variants that receive all citations. We notice that providing citation intents has a negligible effect on faithfulness and decreases the year compliance. This is in line with the overall results. For informativeness, we notice that the same variant wins on all metrics with the exception of insightfulness in the psychology field, where the variant with all citation context and no intents wins by +1%.
G.3 Human-LLM correlation
In Table 12, we report the correlation between LLM and human raters on our evaluation metrics and find that it ranges from moderate to strong, with statistical significance (reported too).
G.4 G-Eval reasoning
We show a few examples of reasons for both high and low scores for each of our metrics in Table 16.
G.5 Visualized impact statements
Examples are shown in Figures 13 (computer science paper),  14 (psychology paper),  15 (psychology paper), and  16 (medicine paper). These impact summaries align closely with our definition. They capture the evolution of citations intents, from initial adoption and theoretical contributions to critiques, methodological refinements, policy influences, and integration into modern applications. For example, the impact summary of “Hierarchically Classifying Documents Using Very Few Words” (ICML’97) illustrates how the paper initially shaped the research paradigm in document classification across fields like bio-informatics and legal document analysis. Later, it faced critiques related to scalability and multi-class accuracy, leading to refining the existing method. Another example is the impact summary of “Sex Bias in Neuroscience and Biomedical Research” (Neuroscience & Biobehavioral Reviews’11), which highlights the early discussions on gender disparities that this paper sparked. Those discussions later led to policy changes, promoting balanced gender representation in preclinical studies, influencing new research methods across various disciplines.
Appendix H More on the human evaluation
H.1 Participants
Nationalities: 2 German, 2 British, 2 Chinese, 1 American, 1 Czech, 1 Albanian; genders: 5 male, 4 female; research areas: AI, NLP, knowledge graphs, databases and information systems, computational social sciences, psychology and brain sciences.
H.2 Perceived usefulness
Results are in Figure 17. We show the overall results in (a), and results for two subsets of papers, namely papers with the most impact-revealing citations in (b), and papers with the highest number of citations in (c). Both subsets show impact summaries that are rated higher, likely because higher counts of impact-revealing citations and citations in general allow for richer contexts, which provide the model with clearer signals about the paper’s impact, leading to more accurate and insightful summaries.
H.3 Open-feedback
A couple professors found certain summaries too generic, failing to highlight specific strengths and limitations. They also pointed out that summaries based on citations from lesser-known conferences may not fully reflect a paper’s true influence, and the quality of summaries often correlates with the size of impact-revealing citation context (a point further validated in Figure 17 (b)). Additionally, one professor noted that for some papers, it was difficult to assess impact because the work was collaborative, and they were only familiar with certain aspects of the paper’s contributions, while their co-authors might be more familiar with other aspects. Other professors suggested improvements in both content and structure: they recommended incorporating citation counts within the text, which we plan to implement in the future, and enhancing the structure of our semi-structured summaries by adding elements like bullet points to improve readability.
Appendix I More on Practical Applications
I.1 Equal citation count, different citing reasons
We select 10 diverse research topics (e.g., LLMs for code generation, cultural sensitivity in clinical psychology), and analyze 3 papers within each. We ensure that the topic-specific papers have close citation counts. In this use case, we demonstrate how fine-grained citation intent analysis provides a deeper understanding of research impact, moving beyond raw citation counts. Results are in Tables 17 and 18. We observe that some papers’ main impact is inspiring or motivating new work, some are used for their method or data, others are cited to point out remaining challenges and propose improvements, etc. For example, for papers on open-information extraction with citation counts of $\sim$ 200,  Cui et al. (2018) is frequently cited to motivate new work or highlight existing limitations, while  Gashteovski et al. (2017) and  Han et al. (2019) are cited for their method. In commonsense knowledge mining, only one of the 3 papers is often cited for its dataset Hwang et al. (2021), and even though the other 2 papers are cited to highlight limitations, the limitations are different, namely challenges with handling constraints and reasoning for one Davison et al. (2019) , but data noise and limited coverage for the other Speer and Havasi (2013). These examples show that even when citation counts are similar, citation intents reveal the specific citing reasons, shaping the perception of a paper’s true impact and assisting in the impact summary generation task.
I.2 Author-level impact summaries
The focus of this paper is to generate impact summaries for individual papers. However, we would like to briefly show how these paper-level summaries can be used to create author-level summaries. Given the top-cited [^13] Top-10 by citation counts. papers of a given author, we generate impact summaries using our best variant, and then ask LLM to aggregate these summaries and infer the author’s overall impact. The prompt used for generating author-level summaries in Figure 18. We show two samples of such summaries for two senior NLP researchers in Figures 19 and  20. This application case hold great potential for future work, and could be extended to research labs, universities, venues, or countries. Methodologically, it would also be particularly interesting to explore different aggregating strategies to generate these summaries, more specifically whether these kind of impact summaries have better quality by aggregating existing paper-level summaries, or by directly use the citation context about an author’s papers regardless of which paper they came from.
Impact summary Analysis and model’s verdict Proof (partial citations list for brevity)
Paper: The importance of race and ethnic background in biomedical research and clinical practice
[2011-2020] Expansion and debate
The discussion around the provided insights flourished, with broader interdisciplinary citations related to ethics, sociological interpretations, and medical trial designs highlighting its significance.
The impact description discusses the flourishing discussion around insights related to race and ethnic background in biomedical research, emphasizing interdisciplinary citations that highlight its significance in ethics, sociological interpretations, and medical trial designs. Several citations in the provided list address the importance of race and ethnicity in biomedical research and clinical practice, indicating that these discussions are indeed significant and supported by various studies.
Verdict: Verifiable
<Associations Between Vitamin D Receptor Polymorphisms and Susceptibility to Ulcerative Colitis and Crohn’s Disease: A Meta-analysis>: "The importance of ethnic background has been raised in bio-medical research and clinical practice.(35) Genetic variation that predisposes to IBD appears to vary between different ethnic groups."
<Conceptual approaches to the study of health disparities.>: "Some researchers have argued that this correlation between genetically identified classifications and self reported race justifies the continued use of self-reported race and ethnicity as a proxy for genetic differences in epidemiologic studies (11, 71)."
Paper: Myocardial infarction redefined–a consensus document of The Joint European Society of Cardiology/American College of Cardiology Committee for the redefinition of myocardial infarction.
[2006 - 2015] Integration into clinical practice guidelines
During this period, the criteria discussed in the input paper were incorporated into widely used clinical guidelines, affirming its essential role in healthcare practice and education.
The impact description states that the criteria discussed in the paper were incorporated into widely used clinical guidelines, affirming its essential role in healthcare practice and education. The citations provided indicate that the Joint European Society of Cardiology/American College of Cardiology criteria for diagnosing myocardial infarction have been referenced and utilized in various studies and guidelines, confirming their integration into clinical practice. This supports the claim that the criteria have been widely adopted in healthcare settings.
Verdict: Verifiable
<Trace element status in Saudi patients with established atherosclerosis.>: "A diagnosis of a myocardial infarction was made in accordance with Joint European Society of Cardiology/American College of Cardiology Committee criteria [17]."
<Improved treatment and prognosis after acute myocardial infarction in Estonia: cross-sectional study from a high risk country>: "The criteria applied for AMI diagnosis on 2001 and 2007 study populations were based on the consensus document published by the European Society of Cardiology in 2000 [12]."
Table 10: Examples of verified (faithful) impact summaries impact descriptions.
Impact summary Themes to be covered
During this early period following the publication, the method proposed by the paper, such as the Spy technique, received attention by its developers who utilized it in the domain of semi-supervised learning. Researchers particularly explored its applications in areas like text document classification and highlighting limitations in existing learning paradigms. The study’s methodology spurred advancements in classifications approaches dealing with positive and unlabeled data. This phase not only saw adaptation of the methodology into various applied fields like web document classification but also critical analysis and theoretical exploration particularly in precision improvement. Researchers evolved the PU learning methods by optimizing the constraints and extending it into more extensive datasets and alternative models leading towards scaling and efficiency enhancements in text and online information domains. In this recent phase, the Spy technique from the discussed paper continues to be influential, inspiring modern machine learning techniques like adaptive PU-learning and text mining approaches. The bridging of neural methods with earlier methodologies showcases the adaptability and robustness over time. Performance Comparison
Novel Algorithm Development
Limitations Acknowledgment
Method Adaptation and Improvement
Applications of PU Learning
Classification Techniques
Data Handling and Sampling
During these years, the paper served as a cornerstone for advancing the implementation of individualized approaches in clinical decision-making. Several studies cited the paper while exploring personalized medicine’s integration. Focus on translational research and advances. This period saw the application of concepts introduced in the paper in fine-tuning medical treatments and drug therapies for individual patients. The references highlight adaptations in various domains from rheumatology to oncology. The paper’s concepts shaped the research direction towards the future application of precision medicine utilizing technologies like AI and computational advances. Advancements in Precision Medicine
Role of Genetic Variants and Pharmacogenomics
Technological Impact on Personalized Medicine
Challenges and Gaps in Personalized Medicine
Patient Variability and Individualized Treatment
Applications of Personalized Medicine in Specific Fields
Ethical and Regulatory Considerations
The input paper, published in 2020, provided a significant synthesis of knowledge about the public health importance of hookworm infections. Effectiveness of Treatment for Anemia
Public Health Significance of Hookworm Disease
Limitations of Targeting School-Age Children
Need for Studies on Military Infection Burdens
..140 more..
Table 11: Examples of automated coverage results (themes actually covered).
Faithfulness Coverage Insightfulness Trend Awareness Specificity
Spearman 0.668* 0.481** 0.489* 0.535* 0.683*
Kendall-Tau 0.668* 0.443** 0.482* 0.533* 0.676*
F1 agreement 0.83 0.71 0.66 0.86 0.74
Table 12: Human and LLM correlation and agreement on evaluating results (Section 3.3). p-value: $\leq$0.001, *$\leq$0.05.
Impact summary Complying citations Noncomplying citations
[2021 - 2022] supporting economic and societal discussions of cardiovascular diseases
Studies during this period heavily relied on the paper for understanding the economic burden and healthcare strategies towards cardiovascular diseases, often referencing the comprehensive statistics presented.
Title: "Extrinsically Conductive Nanomaterials for Cardiac Tissue Engineering Applications",
Year: 2021,
Intent: "highlighting the significant economic impact of cardiovascular diseases"
Title: "Optimising the treatment of chronic ischemic heart disease by training general practitioners to deliver very brief advice on physical activity (OptiCor): protocol of the systematic development and evaluation of a complex intervention",
Year: 2024,
Intent: "highlighting the significant economic impact of cardiovascular events on the German healthcare system"
[2021 - 2023] guiding therapeutic practices and further clinical recommendation development
Multiple citations within this period reflect that the publication served as a cornerstone source for clinical trials and therapeutic evaluation studies aiming to manage PAD effectively.
Title: "Clinical considerations after endovascular therapy of peripheral artery disease",
Year: 2021,
Intent": "highlighting the significance of managing PAD as a CAD equivalent"
Title: "Ankle Brachial Index: An Easy and First-Choice Screening Marker of Peripheral Artery Disease and Physical Function",
Year: 2024,
Intent: "highlighting the significance of ABI in identifying PAD and associated risks"
[2004 - 2010] Initial adoption of stochastic gradient descent methods
After its publication in 2004, the paper introduced stochastic gradient descent (SGD) as an efficient method for solving large-scale linear prediction problems, enabling significant computational advantages in machine learning applications.
Title: "Fast Stochastic Frank-Wolfe Algorithms for Nonlinear SVMs",
Year: 2010,
Intent: "reporting efficiency of stochastic programming techniques"
Title: "A Robbins-Monro Sequence That Can Exploit Prior Information For Faster Convergence",
Year: 2024,
Intent: "highlighting the widespread application and significance of a widely used algorithm"
Table 13: Examples of evidence citations from compliant and noncompliant years. Note the similarity between the summary impact description and the intent of noncompliant citations (highlighted), which might be the source of the error.
Paper Baseline Ours
Liu, Bing, et al. Partially Supervised Classification of Text Documents. 2002. ..[2005-2012]: Development and application
The approaches defined were adapted and extended in a number of practical applications to fine-tune frameworks for domain-specific text classification tasks ..
..[2008-2012]: Applications & theoretical exploration
The methods were applied to problems beyond text classification, such as bioinformatics, personalized recommendations, and data mining. This phase not only saw adaptation of the methodology but also critical analysis and theoretical exploration particularly in precision improvement ..
Martens, James. Deep learning via Hessian-free optimization. 2010. ..[2015-2020]: Adoption in new areas
The techniques described in the paper facilitated advancements in optimizing models in diverse fields like natural language processing and reinforcement learning ..
..[2016-2020]: Integrating into broader optimization frameworks and comparisons
As the field evolved, the approaches from this work became integrated into broader optimization discussions and comparison studies. Researchers began leveraging data-driven adaptations and contextualizing the paper’s contributions relative to emerging advancements like Adam and other methods, enriching overall optimization frameworks ..
Lee, Honglak, et al. Convolutional deep belief networks for scalable unsupervised learning of hierarchical representations. 2009. ..[2013-2017]: Facilitating Advances in Image Recognition
Techniques and methodologies from the paper influenced deep learning architectures, becoming foundational in computer vision applications ..
..[2012-2017]: Advancements in convolutional neural network methodologies
This period marks significant enhancements in deep convolutional techniques, where the paper provided foundational principles that inspired advancements in methodologies across image processing, feature extraction, and specific application areas such as bioinformatics and health diagnostics ..
Myin-Germeys, Inez, et al. Experience sampling research in psychopathology: opening the black box of daily life. 2009. ..[2015-2023]: Application in diverse psychopathological studies
The methodologies introduced by the paper were applied extensively in studies of mood disorders and other psychopathological conditions to dissect complex behavioral and cognitive interactions ..
..[2015-2019]: Expanding applications
The method presented catalyzed its application across diverse practices such as genetic interaction studies, psychological intervention assessments, and computational behavioral analysis. Researchers recognized its role in enhancing ecological validity and precision of dynamic mental state evaluations
[2019-Present]: Technological Integration
This period saw a surge in mobile and wearable technologies for real-time data collection, integrating ESM with machine learning and mobile health platforms to advance personalized medicine and therapeutic interventions. ..
Table 14: Examples of impact summaries from the informativeness evaluation, baseline = zero-shot variant, our method = impact-revealing variants.
Prompt variant Trustworthiness Informativeness
Citations Intents Faith. Cov. Cyc. Insi. Trend. Spec.
Medicine
None 0.70 0.23 n/a 0.75 0.93 0.82
All 0.80 0.30 0.52 0.82 0.95 0.87
All 0.82 0.30 0.47 0.83 0.94 0.88
Impact-rev. 0.85 0.31 0.57 0.82 0.95 0.86
Impact-rev. 0.85 0.32 0.56 0.87 0.96 0.89
Psychology
None 0.74 0.23 n/a 0.75 0.95 0.80
All 0.86 0.32 0.54 0.84 0.98 0.89
All 0.86 0.32 0.52 0.79 0.94 0.86
Impact-rev. 0.90 0.34 0.63 0.82 0.97 0.89
Impact-rev. 0.89 0.32 0.59 0.83 0.98 0.89
Computer science
None 0.87 0.30 n/a 0.60 0.94 0.63
All 0.83 0.35 0.59 0.75 0.96 0.83
All 0.84 0.36 0.46 0.77 0.98 0.82
Impact-rev. 0.85 0.35 0.58 0.74 0.95 0.86
Impact-rev. 0.89 0.37 0.52 0.79 0.99 0.86
Table 15: Ablation results per field - Faithfulness: Faith., Coverage: Cov., Citation Year Compliance: Cyc., Specificity: Spec., Insightfulness: Insi., Trend Awareness: Trend.
Criteria Score Reason
Insightfulness 0.9 The output provides a clear description of the paper’s influence, detailing its foundational, methodological, and expanded relevance. It highlights specific uses and integration into modern frameworks. Slightly more detailed examples could enhance insight further.
Insightfulness 0.4 While the output describes stages of adoption and application, it lacks detailed insight into specific ways the paper directly influenced or was used by other works.
Trend Awareness 1.0 The output clearly identifies how the paper’s impact has changed over time with descriptive period titles. These titles are diverse and informative, fulfilling the evaluation criteria.
Trend Awareness 0.4 The output mentions distinct time periods with descriptions of the paper’s impact, but the titles for the periods are not diverse, as both cover similar themes without clear differentiation.
Specificity 0.9 The output provides detailed periods showcasing the paper’s influence on foundational discussions, specific applications, and technological progress, meeting most evaluation criteria. A minor deduction is due to the lack of explicit mention of individual techniques influenced by the paper.
Specificity 0.3 The output generically states the paper’s influence and provides broad areas of impact, but it lacks concrete details, specific techniques, studies, or frameworks influenced by the paper.
Table 16: Examples of scores and reasons from the LLM evaluator.
Figure 13: Impact of “Hierarchically Classifying Documents Using Very Few Words”, published at ICML’97. (Full example from Figure 1)
Figure 14: Impact of “Successful Adolescent Development among Youth in High-Risk Settings”, published in American Psychologist’93.
Figure 15: Impact of “Comparability of telephone and face-to-face interviews in assessing axis I and II disorders”, published in The American Journal of Psychiatry’97.
Figure 16: Impact of “Sex Bias in Neuroscience and Biomedical Research”, published in Neuroscience & Biobehavioral Reviews’11.
Figure 17: Perceived usefulness results for: (a) all papers, (b) papers in the top 10% for number of impact-revealing citations, (c) papers in the top 10% for total number of citations.
Topic: Open information extraction; citation count: $\sim$200
Paper: Cui et al. (2018) Gashteovski et al. (2017) Han et al. (2019)
Top citation intents: motivating new work in neural OIE, acknowledging limitations in generating facts and schema strength comparing proposed methods with sota approach, method use for OIE method use and extension for relation extraction, highlighting performance issues in capturing various linguistic cues
Topic: Commonsense knowledge mining; citation count: $\sim$350
Davison et al. (2019) Speer and Havasi (2013) Hwang et al. (2021)
acknowledging limitations in handling constraints and reasoning, promoting the perspective of LLMs as knowledge base highlighting limitations of data noise and limited relation coverage, emphasizing importance of commonsense knowledge graphs in AI applications reporting performance comparison and improvements, use as data source for commonsense reasoning
Topic: Retrieval augmented generation; citation count: $\sim$[400-500]
Asai et al. (2023) Ram et al. (2023) Mallen et al. (2023)
motivating the effectiveness of RAG in addressing hallucinations in LLMs, reporting variability in LLMs’ performance with RAG methods motivating the need for well-chosen context in RAG, reporting limitations of retrievers in RAG motivating research on LLMs for question answering, motivating the goal of improving LLMs’ factuality
Topic: LLM as a judge; citation count: $\sim$[370-470]
Fu et al. (2024) Chan et al. (2023) Wang et al. (2024)
motivating the growing interest in using LLMs for automatic evaluations, used on evaluating conversational agents and summary faithfulness highlighting the growing trend of using LLMs for evaluation, use of agent-based methods in evaluation processes motivating the need for further investigation into biases in Judge LLMs, highlighting concerns regarding fairness and bias in using LLMs for evaluation
Topic: Medical domain LLMs; citation count: $\sim$[1900-2100]
Gu et al. (2021) Singhal et al. (2023) Thirunavukarasu et al. (2023)
use of domain-specific pre-trained models, comparing performance of various transformer models in biomedical tasks highlighting a gap in existing research on interactive medical services, performance comparisons of LLMs in medical knowledge tasks highlighting privacy concerns in data collection, performance comparison between LLMs and fine-tuned small models
Topic: Cultural sensitivity in clinical psychology; citation count: $\sim$[1300-1500]
Bernal et al. (2009) Sue (2001) Sue (1998)
highlighting the ongoing debate regarding cultural competency in treatment approaches, motivating the need for comparative analysis of culturally adapted evidence-based practices highlighting limitations in existing multicultural discourse research, motivating the application of cultural competency models in transgender care motivating work on cultural sensitivity in therapy, highlighting the gap in research on treatment efficacy for diverse ethnic populations
Table 17: Similar citation counts, different citation intents. Part I.
Topic: LMs for code generation; citation count: $\sim$[200-300]
Paper: Sun et al. (2020) Le et al. (2022) Mastropaolo et al. (2021)
Top citation intents: drawing inspiration from prior work, comparing various state-of-the-art methods for code generation suggesting potential integration of various search strategies for code generation, motivating the trend of using deep learning for code generation highlighting limitations in existing methods, building on previous ideas for code summarization
Topic: Automated fact checking; citation count: $\sim$700
Ma et al. (2018) Monti et al. (2019) Bian et al. (2020)
motivating the growing interest in ML and NLP for rumor and fake news detection, comparing the proposed model with existing methods for effectiveness verification reporting shortcomings in fake news detection methods, comparing performance improvements of different methods highlighting the limitations of existing datasets for the task, drawing inspiration from existing strategies for early detection and localization
Topic: Multi-modal deep learning; citation count: $\sim$2000
Kim et al. (2021) Liu and Tuzel (2016) Li et al. (2021)
reporting performance comparisons among multimodal methods, highlighting advancements in multimodal pre-training and proposing new tasks reporting model performance compared to state-of-the-art, highlighting the success of GANs in image generation highlighting the limitations of the current method in comparison to other visual LMs, drawing inspiration from existing multimodal models to propose a new method
Topic: Antibiotic resistance mechanisms; citation count: $\sim$[2300-3000]
Aslam et al. (2018) Pang et al. (2019) Wang et al. (2024)
motivating the urgent need for new antibiotics due to rising resistance, highlighting the potential therapeutic effects of the method highlighting the challenges and variations in antibiotic effectiveness against biofilms, motivating the urgent need for novel treatment strategies motivating the need for ongoing antibiotic development and resistance study, highlighting the role of human behavior in antibiotic resistance
Table 18: Similar citation counts, different citation intents. Part II.
Generating a scientific impact summary about a researcher The scientific impact of a researcher can be inferred from the impact of their publications and how they have been used by others. Given a list of impact summaries about papers of a certain researcher, summarize their overall impact, and its evolution over time. This is the list of semi-structured paper summaries about the researcher: \$paper_impact_summaries\$ Generate an impact summary that describes the overall impact of their papers. Focus more on how their papers have been used than what the content of their papers was.
Figure 18: Prompt for generating author-level scientific impact summaries.
Figure 19: The impact summary of the work of a computer science researcher (focus on NLP, cognitive science), H-index=69.
Figure 20: The impact summary of the work of a computer science researcher (focus on NLP, computational linguistics), H-index=61.